OpenStudy (anonymous):

Some birds and squirrels were on a branch of a tree. There were 15 animals on the branch with a total of 50 animal feet. How many birds and how many squirrels were on the branch? Set up the problem like substitution ect. Explain why you set it up that way and solve please

OpenStudy (swissgirl):

Lets start off saying that there are x amount of squirrels. We also know that the total amount of animals is 15 so there wld be 15 minus the amount of squirrels= amt of birds Squirrels = x Birds = 15 -x

OpenStudy (swissgirl):

We also know that there are 50 feet in the branch Squirrels have 4 feet and birds have 2 feet Since we know the total amount of feet we can set up an equation according to the feet.

OpenStudy (swissgirl):

4(x)+2(15-x)=50 4x+30-2x=50 2x+30=50 2x=20 x=10 So we know that there are 10 squirrels on the branch.

OpenStudy (swissgirl):

since we know how many squirrels there are we can figure out how many birds there are Birds: 15-(10)=5 so there are 5 birds one the branch

OpenStudy (anonymous):

ty

OpenStudy (swissgirl):

did u follow i think i rambled on too much hahah